OpenStudy (anonymous):

a=1/2h(b1+b2) solve for b2. How on earth do I do this?

OpenStudy (radar):

That appears to be one of those formulas for area. Isolate b2 using legal algebraic steps.

OpenStudy (anonymous):

But how?

OpenStudy (anonymous):

It matters whether you are talking about 1/2 (h) or 1/(2h). Whichis it?

OpenStudy (anonymous):

The first one. 1/2 is a fraction.

OpenStudy (radar):

There is many ways to go about isolating b2. Here is the path I would take. You do know that multiplying both sides of the equal sign by 2 would still maintain the equality. 2a= h(b1+b2) Did you follow that?

OpenStudy (anonymous):

Yes, I understand that. But do I have to distribute now?

OpenStudy (radar):

Now divide both sides by h. getting (2a)/h = b1 + b2 Still with me?

OpenStudy (anonymous):

Yes.. now do i just subtract b1?

OpenStudy (radar):

I bet you can take it from there.

OpenStudy (radar):

Yes, and be sure what you take away from one side take away from the other.

OpenStudy (anonymous):

so its 2a/h -b1=b2?

OpenStudy (radar):

As I said you could of worked it many different ways, including distributing the h etc.

OpenStudy (radar):

I see you remembered, good job.

OpenStudy (anonymous):

Alright, thank you so much!

OpenStudy (radar):

No problem and you are welcome.